# PhishAgent: A Robust Multimodal Agent for Phishing Webpage Detection

### v1

@@ Abstract
Phishing attacks are a major threat to online security, exploiting user vulnerabilities to steal sensitive information. Various methods have been developed to counteract phishing, each with varying levels of accuracy, but they also encounter notable limitations. In this study, we introduce PhishAgent, a multimodal agent that combines a wide range of tools, integrating both online and offline knowledge bases with Multimodal Large Language Models (MLLMs). This combination leads to broader brand coverage, which enhances brand recognition and recall. Furthermore, we propose a multimodal information retrieval framework designed to extract the top $k$ relevant items from offline knowledge bases, utilizing all available information from a webpage, including logos, HTML, and URLs. Our empirical results, based on three real-world datasets, demonstrate that the proposed framework significantly enhances detection accuracy and reduces both false positives and false negatives, while maintaining model efficiency. Additionally, PhishAgent shows strong resilience against various types of adversarial attacks.
@@ 1 Introduction
To solve these challenges, we propose PhishAgent, a multimodal agent specifically designed for phishing webpage detection with low latency. PhishAgent integrates a comprehensive set of tools, combining both online and offline knowledge bases with MLLMs and independently operating based on a designed workflow. GEPAgent (Wang and Hooi 2024) require multiple iterations of interaction between the agent and knowledge bases to refine results, leading to high latency. In contrast, PhishAgent is designed to use only one such iteration, yet still achieves high performance, thus significantly reducing detection time. Its various modules are interconnected through the Agent Core, which functions as the central module, integrating all the other modules and making decisions. Particularly, the offline knowledge base can cover local brands that are not indexed by search engines, while the online knowledge base can cover very new webpages that the offline knowledge may not include. This combination results in wider brand coverage, thereby increasing brand recognition and recall. Moreover, MLLMs in PhishAgent make decisions by referring to the information queried from the knowledge base rather than solely based on its internal knowledge, which can help enhance the reliability of the results.
@@ 2 Related Works
Reference-based Approaches These compare target webpage information to a known set of brand information. They create a brand knowledge base (BKB) containing logos, aliases, and legitimate domains, and a detector backbone that uses this BKB for detection (Li et al. 2024; Liu et al. 2023). To determine whether a webpage is phishing or legitimate, these systems first identify the target brand of the webpage. If the webpage is found to have the intent of a particular brand but its domain does not align with the brand’s authentic domains, it is classified as phishing. However, these approaches can become outdated and struggle to cover all possible brands comprehensively (Fu, Wenyin, and Deng 2006; Afroz and Greenstadt 2011b; Abdelnabi, Krombholz, and Fritz 2020; Lin et al. 2021; Liu et al. 2022, 2023; Li et al. 2024).
@@
, characterized by LLMs working in conjunction with a set of tools, are especially effective at tackling complex tasks due to their capacity to process information and make informed decisions. Park et al. (2023) introduces generative agents using large language models to simulate human behavior. Works like HuggingGPT (Shen et al. 2024), AgentFLAN (Chen et al. 2024), AgentInstruct (Zeng et al. 2023), and ReAct (Yao et al. 2022) demonstrate the ability of LLMs to manage AI models and solve complex tasks. Recently, GEPAgent (Wang and Hooi 2024), an agent-based phishing detection method, was introduced. It demonstrated high accuracy in identifying phishing webpages by leveraging interactions between MLLMs and an online knowledge base over several iterations. However, it also faced a major drawback regarding execution time, taking an average of 10 seconds per sample.
@@ 3 Threat Model
In a phishing attack, the attacker seeks to deceive users into thinking that the webpage is affiliated with a legitimate brand, thereby tricking them into disclosing sensitive information such as usernames, passwords, or bank details. Formally, let $w$ denote a webpage, which includes its screenshot ($w.$scr), HTML structure ($w.$html), and a URL ($w.$url). To effectively carry out this deception, the webpage must convincingly imitate a specific brand $b$ by leveraging visual elements in $w.$scr, textual features in $w.$html, or both. Our objective is to detect these phishing webpages, correctly identify the intended brand target, and generate detailed annotations that explain why the webpage was classified as a phishing attempt.
@@ 4.1 Overview
We next introduce PhishAgent, which leverages multiple tools and knowledge bases to verify different indicators of phishing activities. The Fig. 1 shows the overview of PhishAgent. PhishAgent consists of four main modules: Preprocessing Module, Online Knowledge-Based Module, Offline Knowledge-Based Module, and Agent Core. Each module can interact with others. The Agent Core functions as the central module, responsible for analyzing information and making decisions, while the remaining modules serve as tools to assist the Agent Core in preprocessing and gathering information. We analyze the details and role of each module in the sections below.
@@ 4.3 Offline Knowledge-Based Module
Figure 2: Our multimodal retriever; Left: Example of how a webpage $w$ and a brand $b$ are encoded and how the retrieval score $s$ is computed; Top right: Example of how the top $k$ brands are retrieved for the webpage $w$ during inference; Bottom right: Example of our training process where contrastive learning is used to distinguish the positive brand (colored in orange) from $N$ randomly sampled negative brands (colored in blue) for the webpage $w$.
@@ 4.5 Agent Core
Agent Core plays a central role in analyzing data returned from the Online and Offline Knowledge-based Modules and making decisions. Agent Core consists of four main components, executed in the following order: Multimodal Brand Extractor, Domain Checker, Target Brand Checker, and Recheck Procedure. The illustration of Agent Core is in Fig. 1. Due to space limitations, prompts and implementation details for each component can be found in the Appendix.
@@
Note that in the case the IBE has been invoked in MBE, the input webpage is directly determined as benign without calling the Recheck Procedure. A case study demonstrating the effectiveness of the Recheck Procedure can be found in the Appendix.
@@ 5 Experiments
   w/o Image-based Brand Extractor 85.88 94.13 95.67 75.15 1.31s
@@ 5.1 Datasets
We utilize three datasets for our main phishing detection experiments. TR-OP (Li et al. 2024) is a manually labeled and balanced dataset, with benign samples randomly selected from the top 50k domains on Tranco (Pochat et al. 2019), and phishing samples sourced from OpenPhish (OpenPhish ). The phishing samples were collected and validated over a six-month period, from July to December 2023, encompassing 440 unique phishing targets. The TR-OP dataset contains 4,000 samples, evenly split between 2,000 phishing and 2,000 benign webpages. SG-SCAN is collected from Singapore’s local webpage traffic. SG-SCAN contains 10k webpages collected from mid-August 2023 to mid January 2024 (Li et al. 2024). SG-SCAN is divided into two datasets. The first dataset, SG-SCAN-1k, is manually labeled and balanced, containing 1,000 webpages, evenly split between 500 phishing webpages and 500 benign webpages. The second dataset, SG-SCAN-unl, is unlabeled and used for the field study. SG-SCAN is used to evaluate the phishing detection approaches in the local context. TR-AP is similar to TR-OP where its benign samples are a different subset of the Tranco top 50k domains from TR-OP. Its phishing samples were gathered from the empirical study of Li et al. (2024), originally from APWG (Anti-Phishing Working Group ). The TR-AP dataset contains 6,000 samples, evenly split between 3,000 phishing and 3,000 benign webpages. The statistics of the 3 datasets can be found in Appendix.
@@ 5.2 Baselines
We enlist three cutting-edge approaches as the phishing detector backbones: Phishpedia (Lin et al. 2021), PhishIntention (Liu et al. 2022), and KPD (Li et al. 2024). As for the knowledge base, Phishpedia and PhishIntention can be integrated with either their original reference list (which includes 277 brands), DynaPhish (increasingly constructed by search engines) (Liu et al. 2023), or KnowPhish (Li et al. 2024). Meanwhile, KPD, due to its requirement for alias information, will be exclusively paired with KnowPhish or DynaPhish, where the extracted brand name is used as the only alias for each new brand. These are reference-based and search engine-based approaches. Additionally, we consider GEPAgent (Wang and Hooi 2024), an agent-based approach, and ChatPhishDetector (Koide et al. 2023), an MLLM-based approach (using ChatGPT 4o), as baselines in our experiments.
@@ 5.5 RQ3: Adversarial Robustness
We investigate the robustness of PhishAgent by simulating various adversarial attacks that could occur in real-world scenarios. These attacks aim to mislead the model into classifying phishing webpages as benign. In practice, attackers can fully manipulate HTML, making it difficult for users to detect the changes. Visual-based attacks are less common because they are more easily noticed by users. We perform these attacks on both HTML and screenshots. The types of adversarial attacks include:
  • Brand spoofing: This attack aims to trick the model into incorrectly identifying a brand name that is not found in the Online Knowledge Base or Offline Knowledge Base, potentially causing the sample to be mistakenly classified as benign based on verification rules. To simulate the attack, we inject adversarial prompts into the HTML, the screenshot (SC), and both.
  • Brand obfuscation: This attack aims to prevent the model from recognizing the brand name. The model may output phrases such as “Not identifiable,” “Does not exist,” or similar phrases. For this attack, we inject adversarial prompts into the HTML or use typosquatting.
We simulate the adversarial attacks on the 500 phishing samples from the SG-SCAN-1k. Detailed implementation of all adversarial attacks can be found in the Appendix.
Table 3 shows the recall of PhishAgent, as well as its performance after removing the Recheck Procedure on different types of attacks. PhishAgent demonstrates robustness against various types of adversarial attacks. Although performance is affected, it remains within acceptable levels, with the greatest reduction being nearly 4% from the original performance when conducting brand spoofing attacks on both screenshots and HTML. Additionally, the experimental results highlight the effectiveness of the Recheck Procedure in enhancing the PhishAgent’s robustness against adversarial attacks.
@@ 5.6 RQ4: Ablation Study
To check the effectiveness of each component in PhishAgent on the overall performance, we sequentially remove each component and observe the changes in performance. We conduct an ablation study on the TR-OP dataset. The accuracy, F1 score, precision, recall, and average running time per sample are reported in Table 4. Overall, all components contribute positively to the phishing detection performance. Further discussion for ablation study experiments can be found in Appendix.

### v2

@@ Abstract
Phishing attacks are a major threat to online security, exploiting user vulnerabilities to steal sensitive information. Various methods have been developed to counteract phishing, each with varying levels of accuracy, but they also face notable limitations. In this study, we introduce PhishAgent, a multimodal agent that combines a wide range of tools, integrating both online and offline knowledge bases with Multimodal Large Language Models (MLLMs). This combination leads to broader brand coverage, which enhances brand recognition and recall. Furthermore, we propose a multimodal information retrieval framework designed to extract the relevant top $k$ items from offline knowledge bases, using available information from a webpage, including logos and HTML. Our empirical results, based on three real-world datasets, demonstrate that the proposed framework significantly enhances detection accuracy and reduces both false positives and false negatives, while maintaining model efficiency. Additionally, PhishAgent shows strong resilience against various types of adversarial attacks.
@@ 1 Introduction
To solve these challenges, we propose PhishAgent, a multimodal agent specifically designed for phishing webpage detection with low latency. PhishAgent integrates a comprehensive set of tools, combining both online and offline knowledge bases. GEPAgent (Wang and Hooi 2024) require multiple iterations of interaction between the agent and knowledge bases to refine results, leading to high latency. In contrast, PhishAgent is designed to use only one such iteration, yet still achieves high performance, thus significantly reducing detection time. Its various modules are interconnected through the Agent Core, which functions as the central module, integrating all the other modules and making decisions. Particularly, the offline knowledge base can cover local brands that are not indexed by search engines, while the online knowledge base can cover very new webpages that the offline knowledge may not include. This combination results in wider brand coverage, thereby increasing brand recognition and recall. Moreover, MLLMs in PhishAgent make decisions based on the information queried from the knowledge base rather than solely based on its internal knowledge, which enhances result reliability.
@@ 2 Related Works
Reference-based Approaches
These compare target webpage information to a known set of brand information. They create a brand knowledge base (BKB) containing logos, aliases, and legitimate domains, and a detector backbone that uses this BKB for detection (Li et al. 2024; Liu et al. 2023). To determine whether a webpage is phishing or legitimate, these systems first identify the target brand of the webpage. If the webpage is found to have the intent of a particular brand but its domain does not align with the brand’s authentic domains, it is classified as phishing. However, these approaches can become outdated and struggle to cover all possible brands comprehensively (Fu, Wenyin, and Deng 2006; Afroz and Greenstadt 2011b; Abdelnabi, Krombholz, and Fritz 2020; Lin et al. 2021; Liu et al. 2022, 2023; Li et al. 2024).
@@
powered by LLMs and tools excel in handling complex tasks through efficient information processing and decision-making. Park et al. (2023) introduces generative agents using large language models to simulate human behavior. Works like HuggingGPT (Shen et al. 2024), AgentFLAN (Chen et al. 2024), AgentInstruct (Zeng et al. 2023), and ReAct (Yao et al. 2022) demonstrate the ability of LLMs to manage AI models and solve complex tasks. Recently, GEPAgent (Wang and Hooi 2024), an agent-based phishing detection method, was introduced. It demonstrated high accuracy in identifying phishing webpages by leveraging interactions between MLLMs and an online knowledge base over several iterations. However, it also faced a major drawback regarding execution time, taking an average of 10 seconds per sample.
@@ 3 Threat Model
In a phishing attack, the attacker seeks to deceive users into thinking that the webpage is affiliated with a legitimate brand, thereby tricking them into disclosing sensitive information such as usernames, passwords, or bank details. Formally, let $w$ denote a webpage, which includes its screenshot ($w.$scr), HTML structure ($w.$html), and a URL ($w.$url). To effectively carry out this deception, the webpage must convincingly imitate a specific brand $b$ by leveraging visual elements in $w.$scr, textual features in $w.$html, or both. Our goal is to detect phishing webpages, identify their target brands, and provide detailed annotations explaination.
@@ 4.1 Overview
We next introduce PhishAgent, which leverages multiple tools and knowledge bases to verify different indicators of phishing activities. The Fig. 1 shows the overview of PhishAgent. PhishAgent consists of four main modules: Preprocessing Module, Online Knowledge-Based Module, Offline Knowledge-Based Module, and Agent Core. Each module can interact with others. The Agent Core functions as the central module, responsible for analyzing information and making decisions, while the remaining modules serve as tools to assist the Agent Core in preprocessing and gathering information. We analyze the details and role of each module in the next sections.
@@ 4.3 Offline Knowledge-Based Module
Figure 2: Our multimodal retriever; Left: Example of how a webpage $w$ and a brand $b$ are encoded and how the retrieval score $s$ is computed; Top right: Example of how the top $k$ brands are retrieved for the webpage $w$ during inference; Bottom right: Example of our training process where contrastive learning is used to distinguish from $N$ randomly sampled negative brands (colored in blue) the positive brand (colored in orange) for the webpage $w$.
@@ 4.5 Agent Core
Agent Core plays a central role in analyzing data returned from the Online and Offline Knowledge-based Modules and making decisions. Agent Core consists of four main components, executed in the following order: Multimodal Brand Extractor, Domain Checker, Target Brand Checker, and Recheck Procedure. The illustration of Agent Core is in Fig. 1. Prompts and implementation details for each component can be found in the Appendix.
@@
Note that in the case the IBE has been invoked in MBE, the input webpage is directly determined as benign without calling the Recheck Procedure.
@@ 5 Experiments
   w/o Image-based Brand Extractor 85.88 84.18 95.67 75.15 1.31s
@@ 5.1 Datasets
We utilize three datasets for our main phishing detection experiments. TR-OP (Li et al. 2024) is a manually labeled and balanced dataset, with benign samples randomly selected from the top 50k domains on Tranco (Pochat et al. 2019), and phishing samples sourced from OpenPhish (OpenPhish ). The phishing samples were collected and validated over a six-month period, from July to December 2023, encompassing 440 unique phishing targets. The TR-OP dataset contains 4,000 samples, evenly split between 2,000 phishing and 2,000 benign webpages. SG-SCAN is collected from Singapore’s local webpage traffic. SG-SCAN contains 10k webpages collected from mid-August 2023 to mid January 2024 (Li et al. 2024). SG-SCAN is divided into two datasets. The first dataset, SG-SCAN-1k, is manually labeled and balanced, containing 1,000 webpages, evenly split between 500 phishing webpages and 500 benign webpages. The second dataset, SG-SCAN-unl, is unlabeled and used for the field study. SG-SCAN is used to evaluate the phishing detection approaches in the local context. TR-AP is similar to TR-OP where its benign samples are a different subset of the Tranco top 50k domains from TR-OP. Its phishing samples were gathered from the empirical study of Li et al. (2024), originally from APWG (Anti-Phishing Working Group ). The TR-AP dataset contains 6,000 samples, evenly split between 3,000 phishing and 3,000 benign webpages.
@@ 5.2 Baselines
We enlist three cutting-edge approaches as the phishing detector backbones: Phishpedia (Lin et al. 2021), PhishIntention (Liu et al. 2022), and KPD (Li et al. 2024). As for the knowledge base, Phishpedia and PhishIntention can be integrated with either their original reference list (which includes 277 brands), DynaPhish (increasingly constructed by search engines) (Liu et al. 2023), or KnowPhish (Li et al. 2024). These are reference-based and search engine-based approaches. Additionally, we consider GEPAgent (Wang and Hooi 2024), an agent-based approach, and ChatPhishDetector (Koide et al. 2023), an MLLM-based approach (using ChatGPT 4o), as baselines in our experiments.
@@ 5.3 RQ1: Effectiveness and Efficiency
PhishAgent mainly classifies webpages as benign when URLs are verified by both knowledge bases. Across benign datasets, 5,236 samples were classified as benign, with only 91 samples (1.74%) due to an unidentifiable brand name, showing knowledge base verification as the primary factor.
PhishAgent operates efficiently in low-resource environments by relying on external APIs for LLMs/MLLMs and online knowledge, minimizing local computation. The retriever requires only  1.6GB of GPU memory, and PhishAgent can function effectively with just online knowledge and MLLMs without the retriever if necessary as shown in RQ4.
@@ 5.5 RQ3: Adversarial Robustness
We evaluate PhishAgent’s robustness against real-world adversarial attacks designed to misclassify phishing webpages as benign. While HTML manipulation is harder for users to detect, visual-based attacks are less common due to their visibility. We test these attacks on both HTML and screenshots (SC). The types of adversarial attacks include:
  • Brand spoofing aims to trick the model into incorrectly identifying a brand name that is not found in the Online Knowledge Base or Offline Knowledge Base, potentially causing the sample to be mistakenly classified as benign based on verification rules. To simulate the attack, we inject adversarial prompts into HTML, SC, and both.
  • Brand obfuscation aims to prevent the model from recognizing the brand name. The model may output phrases such as “Not identifiable,” “Does not exist,” or similar phrases. For this attack, we inject adversarial prompts into the HTML or use typosquatting.
We simulate adversarial attacks on the 500 phishing samples from the SG-SCAN-1k. Table 3 shows the recall of PhishAgent, as well as its performance after removing the Recheck Procedure on different types of attacks. PhishAgent demonstrates robustness against various types of adversarial attacks. Although performance is affected, it remains within acceptable levels, with the greatest reduction being nearly 4% from the original performance when conducting brand spoofing attacks on both screenshots and HTML. Additionally, the experimental results highlight the effectiveness of the Recheck Procedure in enhancing the PhishAgent’s robustness against adversarial attacks.
@@ 5.6 RQ4: Ablation Study
We evaluate the impact of each component in PhishAgent by sequentially removing them and observing performance changes through an ablation study on the TR-OP dataset. Results are shown in Table 4. Overall, all components contribute positively to phishing detection performance. Notably, while removing TBE has minimal impact on accuracy due to IBE’s capabilities, TBE significantly improves efficiency, reducing runtime from 3.35s to 2.25s. Similarly, the Recheck Procedure, although contributing little to overall performance, proves highly effective in adversarial attack scenarios. Further details are provided in the Appendix.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Research Foundation Singapore, NCS Pte. Ltd. and National University of Singapore under the NUS-NCS Joint Laboratory (Grant A-0008542-00-00), as well as by the Ministry of Education, Singapore, through the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00).